OpenStudy (anonymous):

Simplify: (cos^2x-1)/cosx+1

OpenStudy (lalaly):

cos^2x-1= (cosx+1)(cosx-1) so jus cancel out cosx+1 the answer is cosx-1

OpenStudy (anonymous):

top is difference of 2 squares a^2-b^2 = (a+b(a-b)

OpenStudy (anonymous):

man - you guys rock. thank you! never done trig IDs before and all these different formulas are messing me up - I'm forgetting the basics!

OpenStudy (anonymous):

yw - keep practicing - thats so important in maths

OpenStudy (anonymous):

sorry jim - there are no paren. on the numerator, i just did that to show the whole thing was the numerator.....does that change the answer?

OpenStudy (anonymous):

\[\frac{\cos ^2x -1}{\cos x}+1\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no the cosx+1 is the denom.

OpenStudy (anonymous):

(sorry i'm not familiar with the equation editor yet)

OpenStudy (anonymous):

then they did that problem above, layla and Jim

OpenStudy (anonymous):

so cos^2x-1 even without the paren is still the diff. of two squares?

OpenStudy (anonymous):

parenthesis just for organization

OpenStudy (anonymous):

bc cos^2x-1 looks like cosxcosx-1 ... or?

OpenStudy (anonymous):

maybe i should just repost....lol